OpenStudy (anonymous):

What is the value of log81 3? A. y = 4x^2 + 4x + 2 B. y = 2x^2 - 4x + 4 C. y = 4x^2 - 2x - 4 D. y = -2x^2 + 4x - 4

OpenStudy (anonymous):

log (base 81) 3 = log 3/log 81= log 3/ log 3^4= log 3/4log 3= 1/4

OpenStudy (anonymous):

I'm still confused.